Two identicle metal balls with charges +2Q and -Q are separated by some distance and exert a force F on each other .they are now connected by a wire which is then removed .the force between them will now be.??

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Before connecting a wire the force developed is:
F = 2KQ^2/r^2 (attracting)
AFter conecting the wire, the charge gets distributed equally hence,
the charge on each ball becmoes => (2Q – Q)/2 => Q/2.
The new Force b/w them is:
F’ = KQ^2/4r^2
Dividing both of them we get,
F/F’ = 8/1
Henc  ethe new force becomes=> F/8.